# Face Reconstruction Transfer Attack as Out-of-Distribution Generalization

### v1

@@ Abstract
Understanding the vulnerability of face recognition systems to malicious attacks is of critical importance. Previous works have focused on reconstructing face images that can penetrate a targeted verification system. Even in the white-box scenario, however, naively reconstructed images misrepresent the identity information, hence the attacks are easily neutralized once the face system is updated or changed. In this paper, we aim to reconstruct face images which are capable of transferring face attacks on unseen encoders. We term this problem as Face Reconstruction Transfer Attack (FRTA) and show that it can be formulated as an out-of-distribution (OOD) generalization problem. Inspired by its OOD nature, we propose to solve FRTA by Averaged Latent Search and Unsupervised Validation with pseudo target (ALSUV). To strengthen the reconstruction attack on OOD unseen encoders, ALSUV reconstructs the face by searching the latent of amortized generator StyleGAN2 through multiple latent optimization, latent optimization trajectory averaging, and unsupervised validation with a pseudo target. We demonstrate the efficacy and generalization of our method on widely used face datasets, accompanying it with extensive ablation studies and visually, qualitatively, and quantitatively analyses. The source code will be released.
@@
Figure 1: (a) Poorly generalized attack images often get rejected on unseen encoders (middle) while generalized images can bypass other unseen systems (right) like real face images (left).(b) Cosine similarity when only one latent is optimized (red) and the top 1 of when multiple latents are optimized (blue). Red histogram suffers with underfitting. (c) SAR of top 1 to 5 candidates tested on unseen encoders of target encoder (coral bar) and our method (turquoise bar). Our method shows consistent correlation between rank and performance as well as better results.
@@ 1 Introduction
With the increasing deployment of face recognition systems in security-critical environments, threat actors are developing sophisticated attack strategies over various attack points, where one of the major threats is face reconstruction attacks [18, 6, 7, 23, 24, 8]. The primary goal of face reconstruction attacks is to create fake biometric images that resemble genuine ones from the stored biometric templates which are then used to bypass the system. Previous works have mostly focused solely on attacking the target (seen) encoder, i.e., using these fake biometric images to bypass the same system. However, transfer attack scenarios, where these fake biometric images are used to bypass other unseen systems (Fig. 1a middle) are not discussed enough. They are potentially more perilous than common attacks as they can break into a wide range of face recognition systems.
@@
Formally, we define Face Reconstruction Transfer Attacks (FRTA) as successfully reconstructing a face image that can substitute a real face image on unseen encoders, as illustrated in Fig 1a. To state our problem in a rigorous and tractable framework, we formulate this task to reconstruct a face image which matches the original image in identity by a finite number of unseen face encoders given only a single encoder and a feature embedded through this encoder (section 3). However, existing works do not consider transfer attacks in their designs thoroughly.
@@
To solve FRTA effectively, we first devise a novel out-of-distribution (OOD) oriented FRTA framework that reformulates the attack as a problem of generalization of loss function over OOD of network parameters. Our FRTA task falls to the standard OOD generalization category in that the loss function needs to be optimized by one variable and generalized with respect to the other instantiated from unseen distributions. We postulate a white-box scenario where a single encoder and a template feature embedded through this encoder are given and face reconstruction is achieved by optimizing the data input in a way to minimize feature reconstruction loss. Instead of directly updating the input image, we adopt a generative model and update its latent which outputs an image(Fig. 2 step 1).
@@
However, reconstructing the face with naive latent optimization is likely to suffer from underfitting with poor latent optimizations (Fig. 1b red histogram). To address this challenge, we introduce an Averaged Latent Search with Unsupervised Validation through pseudo target (ALSUV) framework, which is motivated by the OOD generalization concept. In this OOD-oriented approach, we 1) optimize multiple latents concurrently by 2) employing latent averaging and 3) searching for the most optimal generalized sample through unsupervised validation using a validation encoder. Multiple latent optimization prevents poor optimization by selecting the well-optimized sample close to the target(Fig. 1b blue histogram). However, this potentially causes overfitting to the seen encoder (Fig. 1c coral bars). Therefore, we average latents throughout optimization trajectories which provides a flatter loss surface (Fig. 4) and leads to OOD generalization [21, 3, 22]. Additionally, we adopt an unsupervised validation search with pseudo target, which incorporates a surrogate validation encoder to search for the best generalizing sample in validation encoder space. Overall, latent averaging and unsupervised validation alleviate overfitting to seen encoder as shown in Fig. 1c (turquoise bar) where rank and performance are notably correlated.
Contributions
The contributions of our work are summarized as:
  • We address FRTA as a threat that can potentially bypass a wide range of unseen face recognition systems. We rigorously formulate the reconstruction problem as an OOD generalization problem and enhance the attack performance with a multiple latents optimization strategy. To address overfitting, we introduce the OOD-oriented ALSUV, which includes latent averaging and unsupervised validation using pseudo targets from a validation encoder.
@@
  • Our method achieves state-of-the-art in face reconstruction transfer attack tested on a subset of LFW, CFP-FP, and AgeDB datasets with 6 different types of face encoders in terms of conventionally used success attack rate and top-1 identification rate where we are the first to attempt.
  • We comprehensively examine our method through ablation study, hyperparameter analysis, and evaluate the image quality. Also, we conduct additional experiments to analyze the effect of optimizing multiple latents, latent averaging, and unsupervised validation.
@@ 2.1 Face Reconstruction from Features
NBNet [18] first pioneered face reconstruction from the template by neighborly deconvolution, but the results are substandard for both quality and performance. [6] projects features into the latent space of a pre-trained StyleGAN2 [14] to generate fine-grained resembling images. The results are qualitatively decent, but often contain different identities as shown in Fig. 6. DiBiGAN [8] presents a generative framework based on bijective metric learning and pairs features with face images one-to-one. These methods offer fast sample generation after training but require extensive face datasets and time for training a new network. [23] samples varying random Gaussian blobs iteratively and combining the blobs as the shape of a face. It requires only a few queries and no prior knowledge such as dataset, but shows results with low quality. [24] reconstructs faces using similarity scores based on eigenface with soft symmetry constraints, generative regularization, and multi-start policy to avoid local minimas. However, the reconstructed images show severe noise as shown in Fig. 6. These works consume less time to generate few samples than training a new network but takes longer for large scale generations.
Methods introduced above show promising results when tested with seen encoders, but performance drastically falls with unseen encoders. Recently, a couple of works [7, 19, 27] considered FRTA scenarios in their work. [7] suggests a genetic algorithm-based approach along with an attack pipeline to impersonate the target user. Reconstructed images are high quality, but evolutionary algorithms rely on random mutation and selection processes which might get trapped in local optima or fail to generalize well due to the limited exploration-exploitation trade-off inherent in their design. [19] suggests query efficient zeroth-order gradient estimation with top $k$ initialization search followed by ensembling. This approach typically iteratively adjusts latent representations to minimize reconstruction errors but may suffer from overfitting to specific characteristics of the seen encoder. [27] trains a network which maps features to the $\mathcal{W^{+}}$ latent space of StyleGAN3 to learn the $\mathcal{W^{+}}$ distribution space based on a WGAN framework, however, the GAN frameworks are prone to mode collapse.
@@ 3 Face Reconstruction Transfer Attack as OOD Generalization
We first clearly define the problem of FRTA, and we show that face reconstruction transfer attack is in fact an OOD generalization problem.
@@ 3.1 Problem Formalization
The FRTA can be formalized as an algorithm $\mathcal{A}$ that must return a solution vector $\mathcal{A}(\theta_{seen})=x^{*}$ to the following optimization objective:
@@
One approach to the attack is by using a face image generator $G$. By substituting $x=G(z)$ in the above equation, one maximizes the objective in terms of $z$ instead of $x$ as:
$\max_{z}\min_{\theta\in\Theta}sim(E_{\theta}(G(z)),v_{\theta}).$ (2)
@@ 3.3.1 Multiple Latent Optimization
In order to avoid the underminimization problem shown in Fig. 1b and to generate candidates for our following unsupervised validation method, we initialize multiple $n$ latent vectors and optimize them in a parallel manner:
@@
where $E_{seen}=E_{\theta_{seen}}$ $v_{seen}=v_{\theta_{seen}}$. The given loss function is minimized by a gradient-based update using the standard optimizer such as Adam[15] or SGD. Fig. 3 validates that updating with multiple latents significantly improves the minimization of the loss. Moreover, we find that this simple multiple latent optimization can more effectively escape from poor local minima than iterating with complicated learning rate scheduler (Tab. 6).
@@ 3.3.2 Latent Averaging
Avoiding the poor underminimization issue alone is not sufficient for effective generalization of the attack under FRTA since the acquired solution may overfit to seen encoders as shown in Fig. 1c. To effectively improve the attack rate on the unseen encoders, we borrow the idea from OOD generalization[21, 3, 22] and apply averaging the solution latent vectors over the optimization trajectory:
$\overline{z}_{i}=\frac{1}{T_{0}}\sum_{t=T-T_{0}}^{T}z^{(t)}_{i}$ (7)
where $z^{(t)}_{i}$ are the latent vectors acquired at step $t$ of the optimization of Eq. (6), $T$ is the total number of optimization steps, and $T_{0}$ is the size of steps to average the latent vectors.
Due to the equivalence between FRTA and OOD generalization (Thm. LABEL:thm:taood), latent averaging can improve the generalization of the similarity maximization under unseen encoder networks and corresponding unseen targets. Fig. 3 evidences our hypothesis, indicating that averaging the latents improves the attack rate on the unseen encoders. Moreover, Fig. 4 shows that latent averaging improves the transfer attack by smoothening the loss surface of our objective, validating the equivalence between OOD generalization and FRTA.
@@ 3.3.3 Unsupervised Validation with Pseudo Target
Despite its effectiveness, multiple latent optimizations with averaging can still suffer overfitting issues as no explicit information of unseen encoders is exposed to the optimization. Therefore, we acquire more explicit information of unseen encoders by utilizing a surrogate validation encoder $E_{val}=E_{\theta_{val}}$. Particularly, we propose to validate our reconstruction criterion Eq. (1) over the surrogate validation encoder. One remaining issue however is the absence of validation target vector $v_{val}$. To resolve it, we construct the pseudo target:
@@ 3.3.4 Full Objective
Overall, the ALSUV algorithm searches the solution latent $z^{*}$ by unsupervised validation
@@ 4 Experiments
The experiments section includes: 1) performance evaluation against existing methods; 2) comprehensive component ablation and hyperparameter variation to show effectiveness; 3) analysis of components by varying setups, comparing parallel latent optimization to serial optimization, visualizing loss surface effects with and without latent averaging, using different validation encoders for unsupervised validation, and assessing image quality visually and quantitatively.
@@ 4.1 Configuration
We use StyleGAN2[14] trained with FFHQ-256 for the generative model, denoted as $G(\cdot)$, and latents are optimized in the $\mathcal{W^{+}}$ space. Both $G(\cdot)$ and target encoders $E_{seen}(\cdot)$ are frozen while optimizing latents. We adopt Adam[15] optimizer with 100 steps where the learning rate starts from 0.1 and is divided by 10 at iteration 50. Our method involves three hyperparameters: $n=100$, the number of latents; $t=70$, length of trajectory for latent averaging; and $k_{top}=10$, the number of samples used for unsupervised validation. We use pytorch[20] for all experiments on a single Nvidia RTX 2080ti GPU.
Table 1: Success Attack Rate (SAR) of previous works and our proposed method. Red colored cells indicate results of seen encoder setup. We highlight our method with grey and best results are highlighted with bold.
@@
Table 2: Rank 1 identification rate of previous works and our proposed method. Red colored cells indicate results of seen encoder setup. We highlight our method with grey and best results are highlighted with bold.
@@ 4.2 Datasets and Networks
We use the LFW, CFP-FP, and AgeDB-30 datasets, three widely used verification datasets with distinct characteristics. For LFW and AgeDB-30, we compare reconstructed samples with every other positive sample except the one used for reconstruction, resulting in 3,166 and 3,307 pairs of comparisons, respectively. CFP-FP consists of frontal and profile images; we reconstruct only the frontal images and follow the challenging frontal-profile verification protocol, resulting in 130 pairs in total (not all sampled identities were used in the given protocol). For identification, we set up generated samples as probes and every image as the gallery, resulting in 13,233, 2,000, and 5,298 samples in the gallery, respectively, consisting of 5,749, 500, and 388 identities for LFW, CFP-FP, and AgeDB-30, respectively. For CFP-FP, we use generated frontal images for probes and profile images for the gallery, which is very challenging. We randomly sample 200 non-overlapping identities each from 3 datasets. We use encoders based on various backbones which are equipped with different classification heads and distinct datasets. Specific configurations are shown in Supp. 0.D Tab. 8. For the validation encoder, we use Swin-T as the default.
@@ 4.3 Evaluation Metrics and Details
[18] introduces Type I and Type II SAR where Type I compares the generated face with the ground truth target, while Type II compares with different images from the same identity. SAR measures the ratio of generated samples passing the positive verification test where thresholds are specific to type of datasets and face encoders. Since Type I is relatively easy, we only report Type II performance which is more challenging. We also evaluate identification rate, where we retrieve the top 1 sample from gallery composed of real samples and probes consisting of generated samples. We include target samples in the gallery which is still challenging. In ablation, we also report the SAR@FAR which is the success attack rate at thresholds of FAR($=1e-4,1e-3,1-e2$).

### v2

@@ Abstract
Understanding the vulnerability of face recognition systems to malicious attacks is of critical importance. Previous works have focused on reconstructing face images that can penetrate a targeted verification system. Even in the white-box scenario, however, naively reconstructed images misrepresent the identity information, hence the attacks are easily neutralized once the face system is updated or changed. In this paper, we aim to reconstruct face images which are capable of transferring face attacks on unseen encoders. We term this problem as Face Reconstruction Transfer Attack (FRTA) and show that it can be formulated as an out-of-distribution (OOD) generalization problem. Inspired by its OOD nature, we propose to solve FRTA by Averaged Latent Search and Unsupervised Validation with pseudo target (ALSUV). To strengthen the reconstruction attack on OOD unseen encoders, ALSUV reconstructs the face by searching the latent of amortized generator StyleGAN2 through multiple latent optimization, latent optimization trajectory averaging, and unsupervised validation with a pseudo target. We demonstrate the efficacy and generalization of our method on widely used face datasets, accompanying it with extensive ablation studies and visually, qualitatively, and quantitatively analyses. Code: https://github.com/jungyg/ALSUV.git
@@
Figure 1: (a) Poorly generalized attack images often get rejected on unseen encoders (middle) while generalized images can bypass other unseen systems (right) like real face images (left).(b) Cosine similarity when only one latent is optimized (red) and the top 1 of multiple latent optimized (blue). Red histogram suffers with underfitting. (c) SAR of top 1 to 5 candidates of target encoder (coral bar) and our method (turquoise bar) tested on unseen encoders. Our method shows consistent correlation between rank and performance as well as better results.
@@ 1 Introduction
With the increasing deployment of face recognition systems in security-critical environments, threat actors are developing sophisticated attack strategies over various attack points, where one of the major threats is face reconstruction attacks [21, 7, 8, 26, 27, 9, 30, 31, 32]. The primary goal of face reconstruction attacks is to create fake biometric images that resemble genuine ones from the stored biometric templates which are then used to bypass the system. Previous works have mostly focused solely on attacking the target (seen) encoder, i.e., using these fake biometric images to bypass the same system. However, transfer attack scenarios, where these fake biometric images are used to bypass other unseen systems (Fig. 1a middle) are not discussed enough. They are potentially more perilous as they can break into a wide range of face recognition systems.
@@
Formally, we define Face Reconstruction Transfer Attacks (FRTA) as successfully reconstructing a face image that can substitute a real face image on unseen encoders, as illustrated in Fig 1a. To state our problem in a rigorous and tractable framework, we formulate this task to reconstruct a face image which matches the original image in identity by a finite number of unseen face encoders given only a single encoder and a feature embedded through this encoder (section 3). However, existing works pay less attention to transfer attacks.
@@
To solve FRTA effectively, we first devise a novel out-of-distribution (OOD) oriented FRTA framework that reformulates the attack as a problem of generalization of loss function over OOD of network parameters. Our FRTA task falls to the standard OOD generalization category in that the loss function needs to be optimized by one variable and generalized with respect to the other instantiated from unseen distributions. We postulate a white-box scenario [^1] The experiment in black-box scenario is given in Sec. 0.C.6 where a single encoder and a template feature embedded through this encoder are given, and face reconstruction is achieved by optimizing the data input in a way to minimize feature reconstruction loss. Instead of directly updating the input image, we adopt a generative model and update its latent which outputs an image (Fig. 2 step 1).
@@
However, reconstructing the face with naive latent optimization is likely to suffer from underfitting with poor latent optimization (Fig. 1b red histogram). To address this challenge, we introduce an Averaged Latent Search with Unsupervised Validation through pseudo target (ALSUV) framework, which is motivated by the OOD generalization concept. In this OOD-oriented approach, we 1) optimize multiple latents concurrently by 2) employing latent averaging and 3) searching for the most optimal generalized sample through unsupervised validation using a validation encoder. Multiple latent optimization prevents poor optimization by selecting the well-optimized sample close to the target (Fig. 1b blue histogram). However, this potentially causes overfitting to the seen encoder (Fig. 1c coral bars). Therefore, we average latents throughout optimization trajectories which provides a flatter loss surface (Fig. 4) and leads to OOD generalization [24, 4, 25]. Additionally, we adopt an unsupervised validation search with pseudo target, which incorporates a surrogate validation encoder to search for the best generalizing sample in validation encoder space. Overall, latent averaging and unsupervised validation alleviate overfitting to seen encoder as shown in Fig. 1c (turquoise bar) where rank and performance are notably correlated. The contributions of our work are summarized as:
  • We rigorously formulate the reconstruction transfer attack as an OOD generalization problem. Inspired by this framing, we effectively improve the transfer attack performance by an OOD-oriented method ALSUV.
@@
  • Our method achieves state-of-the-art in the transfer attack scenarios based on LFW, CFP-FP, and AgeDB with 6 different types of face encoders under the measures of success attack rate and top-1 identification rate.
  • We conduct in-depth empirical analysis on the components of ALSUV. Particularly, we observe that ALSUV improves the transfer attack performance by finding a flatter minima in the reconstruction loss landscape, evidencing the close connection between face transfer attack and OOD generalization.
@@ 2.1 Face Reconstruction from Features
NBNet [21] first pioneered face reconstruction from template by neighborly deconvolution, but the results are substandard for both quality and performance. [7] projects features into the latent space of a pre-trained StyleGAN2 [16] to generate fine-grained resembling images. The results are qualitatively decent, but often contain different identities as shown in Fig. 6. DiBiGAN [9] presents a generative framework based on bijective metric learning and pairs features with face images one-to-one. These methods offer fast sample generation after training but require extensive face datasets and time for training a new network. [26] samples varying random Gaussian blobs iteratively and combining the blobs as the shape of a face. It requires only a few queries and no prior knowledge such as dataset, but shows results with low quality. [27] reconstructs faces using similarity scores based on eigenface with soft symmetry constraints, generative regularization, and multi-start policy to avoid local minimas. However, the reconstructed images show severe noise as shown in Fig. 6. These works consume less time to generate few samples than training a new network but takes longer for large scale generations.
Methods introduced above show promising results when tested with seen encoders, but performance drastically falls with unseen encoders. Recently, [8, 22, 32, 31] regarded transfer attack scenarios in their works. [8] suggests a genetic algorithm-based approach along with an attack pipeline to impersonate the target user. Reconstructed images are high quality, but evolutionary algorithms relying on random mutation and selection processes might get trapped in local optima or fail to generalize well due to the limited exploration-exploitation trade-off in their inherent design. [22] suggests query efficient zeroth-order gradient estimation with ensembling the initial top $k$ initialization search followed by ensembling. This approach iteratively adjusts latent representations to minimize reconstruction errors but suffer from overfitting to specific characteristics of the seen encoder. [32, 31] train networks which map face embeddings to the $\mathcal{W^{+}}$ latent space of StyleGAN3 and 3D aware generative model respectively to utilize the representation of each generative models. They evaluate 5 different types of attacks including transfer attack in whitebox and blackbox scenario in their work. Unlike previous works, we approach FRTA as an OOD generalization problem with more challenging datasets and various types of face recognition systems.
@@ 3 Face Reconstruction Transfer Attack as OOD Generalization
We first rigorously define the problem of FRTA, and we show that face reconstruction transfer attack is in fact an OOD generalization problem.
@@ 3.1 Problem Formalization
The FRTA can be formalized as an algorithm $\mathcal{A}$ that must return a solution vector $\mathcal{A}(\theta_{seen})=x^{*}$ to the optimization objective
@@
One approach to the attack is by using a face image generator $G$. By substituting $x=G(z)$ in the above equation, the objective is maximized in terms of $z$ instead of $x$: $\max_{z}\min_{\theta\in\Theta}sim(E_{\theta}(G(z)),v_{\theta}).$
@@ 3.3.1 Multiple Latent Optimization
In order to avoid the underminimization problem shown in Fig. 1b and generate candidates for our following unsupervised validation method, we initialize multiple $n$ latent vectors and optimize them in a parallel manner:
@@
where $E_{seen}=E_{\theta_{seen}}$ $v_{seen}=v_{\theta_{seen}}$. The given loss function is minimized by a gradient-based update using the standard optimizer such as Adam[18] or SGD. Fig. 3 validates that updating with multiple latents significantly improves the minimization of the loss. Moreover, this method can effectively escape from poor local minima than iterating with complicated learning rate scheduler (Tab. 6).
@@ 3.3.2 Latent Averaging
Avoiding the poor underminimization issue alone is not sufficient for effective generalization of attack under FRTA since the acquired solution may overfit to seen encoders as shown in Fig. 1c. To improve the attack rate on the unseen encoders, we borrow the idea from OOD generalization[24, 4, 25] and apply averaging the solution latent vectors over the optimization trajectory:
$\overline{z}_{i}=\frac{1}{t_{avg}}\sum_{t=T-t_{avg}}^{T}z^{(t)}_{i}$ (6)
where $z^{(t)}_{i}$ are the latent vectors acquired at step $t$ of the optimization of Eq. (5), $T$ is the total number of optimization steps, and $t_{avg}$ is the size of steps to average the latent vectors.
Due to the equivalence between FRTA and OOD generalization from the theorem above, latent averaging can improve the generalization of the similarity maximization under unseen encoder networks and corresponding unseen targets. Fig. 3 evidences our hypothesis, indicating that averaging the latents improves the attack rate on the unseen encoders. Moreover, Fig. 4 shows that latent averaging improves the transfer attack by smoothening the loss surface of our objective, validating the equivalence between OOD generalization and FRTA.
@@ 3.3.3 Unsupervised Validation with Pseudo Target
Despite its effectiveness, multiple latent optimization with averaging can still suffer overfitting issues as no explicit information of unseen encoders is exposed to the optimization. Therefore, we acquire more explicit information of unseen encoders by utilizing a surrogate validation encoder $E_{val}=E_{\theta_{val}}$. Particularly, we propose to validate our reconstruction criterion Eq. (1) over the surrogate validation encoder. One remaining issue however is the absence of validation target vector $v_{val}$. To resolve it, we construct the pseudo target:
@@ 3.3.4 Full Objective
Overall, ALSUV searches the solution latent $z^{*}$ by unsupervised validation
@@ 4 Experiments
The experiments section includes: 1) performance evaluation against existing methods; 2) comprehensive component ablation and hyperparameter variation to show effectiveness; 3) component-wise in-depth ablation study and analysis on ALSUV.
@@ 4.1 Configuration
We use StyleGAN2[16] trained on FFHQ-256 for the generator $G$, and the latents are optimized in the $\mathcal{W^{+}}$ space. Both $G$ and target encoders $E_{seen}$ are frozen while optimizing latents. We adopt Adam[18] optimizer with 100 steps where the learning rate starts from 0.1 and is divided by 10 at iteration 50. Our method involves three hyperparameters: $n=100$, the number of latents; $t_{avg}=70$, length of trajectory for latent averaging; and $k_{top}=10$, the number of samples used for unsupervised validation. We use pytorch[23] for all experiments on a single Nvidia RTX 2080ti GPU.
Table 1: SAR of previous works and our proposed method. Red cells indicate results of seen encoder setup. We highlight our method with grey and best results are highlighted with bold.
@@
Table 2: Rank 1 identification rate of previous works and our proposed method. Red cells indicate results of seen encoder setup. We highlight our method with grey and best results are highlighted with bold.
@@ 4.2 Datasets and Networks
We use LFW, CFP-FP, and AgeDB-30, three widely used verification datasets with distinct characters. We randomly sample 200 non-overlapping identities each from 3 datasets. Encoders are selected with distinct backbones equipped with various classification heads trained with distinct datasets. We use pre-trained models from [37] and models trained by our own trained with CASIA-WebFace[39] or MS-1M[11]. Specific configurations of dataset and encoders are shown in Supp. 0.D.1. For the validation encoder, we use Swin-T as default.
@@ 4.3 Evaluation Metrics and Details
[21] introduces Type I and Type II SAR where Type I compares the generated face with the ground truth target image, while Type II compares with different images with the same identity. SAR measures the ratio of generated samples passing the positive verification test where thresholds are specific to type of datasets and face encoders. Since Type I is relatively easy, we only report Type II performance which is more challenging. We also evaluate identification rate, which retrieves the top 1 sample from a gallery set composed of real images with probes consisting of generated samples. We include target samples in the gallery.
In ablation, we also report the SAR@FAR which is the success attack rate at thresholds of FAR($=1e-4,1e-3,1-e2$).